Soybean meal is 12% protein. Cornmeal I'd 6% protein. How many pounds of each should be mixed to get a 240 pound mixture that is 10% protein?
Answer by ankor@dixie-net.com(22740)

Soybean meal is 12% protein. Cornmeal is 6% protein.
How many pounds of each should be mixed to get a 240 pound mixture that is 10% protein?
:
Let x = amt of 12% protein
then
(240-x) = amt of 6% protein
:
At typical mixture equation:
.12x + .06(240 - x) = .10(240)
.12x + 14.4 - .06x = 24
.12x - .06x = 24 - 14.4
.06x = 24 - 14.4
.06x = 9.6
x = 9.6/.06
x = 160 lb of soy meal
and
240 - 160 = 80 lb of corn meal
:
:
Check solution
.12(160) + .06(80) = .10(240)